# Transport evidence of surface states in magnetic topological insulator MnBi2Te4

### v1

@@ Abstract
One convenient way to individually access the doping and energy levels of the surface and bulk electronic populations is to perform Landau level spectroscopy. The measurement of Shubnikov-de-Haas oscillations (SdHO) of the resistance under external magnetic field can give access to the carrier density, the transport and the quantum mobilities, the effective mass and the chemical potential of different charge carrier populations, whose contributions to the total conductance can be disentangled so. However, this technique requires magnetic fields such that $\mu B>1$, $\mu$ being the charge carrier mobility, different for each band. While such quantum oscillations have been studied in Sb-doped MnBi2Te4,[Jiang2021, Lee2021], they were not observed in pure MnBi2Te4 in the many magneto-transport studies up to 15 T[Cui2019, Lee2019, Chen2019a, Bac2022, Martini2023, Cao2024], except for the report of small resistance oscillations [Lei2022]. This absence of SdHO in pure MnBi2Te4 is most probably due to a low mobility caused by the strong disorder in the bulk, attributed mainly to Mn/Bi antisites [Otrokov2019, Yan2019, Liu2020, Li2024b, Souchay2019, Zeugner2019], thus requiring very high magnetic fields to observe SdHO.
Figure 1: Device characterization. a: Atomic force microscopy picture of an etched MnBi2Te4 nanostructure, with thickness 91 nm and roughness of about 1 nm. The original shape of the MnBi2Te4 exfoliated flake can still be faintly seen below the contacts. b: temperature dependence of the longitudinal resistance in zero magnetic field. The antiferromagnetic transition induces a resistance peak at 25.5 K.
@@
Figure 2: High-field magnetotransport in a MnBi2Te4 nanostructure at 4.2 K. a: Magnetoresistance and b: Hall effect up to 55 T. The magnetoresistance and Hall effect have been respectively symmetrized and antisymmetrized with magnetic field. Shubnikov-de-Haas oscillations are visible above about 40 T. Purple dashed lines represent respectively a cubic fit of the magnetoresistance, used afterward to extract Shubnikov-de-Haas oscillations, and a linear fit of the Hall effect asymptote at high field. Arrows indicate the spin-flop- and saturation fields.
Fig.2 presents the longitudinal and the transverse magnetoresistance under perpendicular magnetic fields up to 55 T. At low field, both samples show the expected signature of a collinear antiferromagnet with a uniaxial anisotropy: a sharp resistance peak at about 3.1 T corresponds to the spin-flop magnetic transition,[Tan2020] which is also visible in the transverse resistance as a sharp jump of the anomalous Hall effect (blue arrows). Another feature is observed at 7 T, corresponding to the saturation field above which the magnetization is uniform (see Supplementary Information). Above 10 T, the transverse resistance is linear up to 55 T, corresponding to an asymptote of the normal Hall effect, yielding a total bulk carrier density of $6.7\times 10^{19}$ cm-3.
Figure 3: Temperature dependence of Shubnikov-de-Haas oscillations. a: Residual Shubnikov-de-Haas oscillations after removal of a cubic background, versus inverse magnetic field. The dashed lines show the position of the SdHO extrema, separated by 0.003 T-1. b: Residual Shubnikov-de-Haas oscillations after removal of a cubic background, at different temperatures. The position of the minimum and maximum of the oscillation are indicated by dashed lines. c: Temperature dependence of the Shubnikov-de-Haas oscillation amplitude for the peaks at 43 and 49T. The error bars correspond to rapid noise in the raw data. Lifshitz-Kosevich fits lead effective masses $m^{*}$ of about $0.15-0.17m_{\mathrm{e}}$.
@@
In order to determine the 3D or 2D character of the electronic population responsible for the observed SdHO, the magnetic field was tilted out of the sample plane by an angle $\theta$. From the angular dependence of the magnetoresistance we observe that the SdHO evolves with the angle, but is stable with respect to $B_{\perp}=B\cos(\theta)$. Fig.4a shows the SdHO plotted against the perpendicular component of the magnetic field $B_{\perp}$. The position of the maximum and minimum of the visible SdHO remains unaffected, while the positions vary with regard to the total field. This is a clear sign that the SdHO originate from a 2D electronic state in the sample plane. This is also compatible with the carrier density extracted from the SdHO frequency: if interpreted as a 3D parabolic bulk band, a magnetic frequency $f_{B}=167$ T corresponds to a bulk carrier density of $n_{\mathrm{3D}}^{\mathrm{SdHO}}=1/2\pi^{2}\,(2e/h\times f_{\mathrm{B}})^{3/2}=8\times 10^{17}$ cm-3, which is two orders of magnitude smaller than the total carrier density extracted from the Hall effect. Interpreting the SdHO as originating from a topological surface state (TSS) with linear energy dispersion and a Fermi velocity $v_{\mathrm{F}}=5.5\times 10^{5}$ m.s-1,[Tomarchio2022, Gong2019] we calculate a 2D carrier density of $n_{\mathrm{2D}}^{\mathrm{SdHO}}=e/h\times f_{\mathrm{B}}=4.1\times 10^{12}$ cm-2 at the topological surface state, and a Fermi energy 255 meV above the Dirac point in good agreement with ARPES [Xu2024, Estyunin2020].
@@
Figure 4: Angular dependence of Shubnikov-de-Haas oscillations and band bending. a: Angular dependence of the Shubnikov-de-Haas oscillations after removal of a cubic background, between 0°  (transverse field configuration) to 30°  tilt toward in-plane configuration, shown against the transverse magnetic field component $B_{\perp}$. Curves shifted for clarity. The schematic shows the tilting angle $\theta$ of the magnetic field. b: Schematic of the energy levels and band bending in the MnBi2Te4 nanostructure, as extracted from Shubnikov-de-Haas and Hall data (see main text), in a model with three conduction bands (CB1, CB2, CB3, see main text) and one topological surface state (TSS).
@@
According to ARPES measurements on MnBi2Te4, the Dirac point lies about 100 meV below the bottom of the lowest conduction band CB1,[Li2019f, Hao2019, Estyunin2020], which, together with the $E_{\mathrm{F}}$ of 255 meV above the Dirac point as calculated above, results in an upward band bending of more than 150 meV between bulk (CB1 minimum at 317 meV) and surface (CB1 minimum at 155 meV). This upward band bending is a common situation in heavily doped 3DTI [Veyrat2015], where topological surface states are filled by charge transfer from heavily doped bulk bands, resulting in a lower doping of the bulk band close to the topological surface over the length-scale of the Thomas-Fermi screening length. It is, however, interesting to notice that, given the important band bending between TSS and bulk states, surface-sensitive techniques such as ARPES measurements would significantly underestimate the bulk chemical potential, and in particular the filling of the third conduction band CB3. Given the large effective mass of CB3, the bulk chemical potential will efficiently be pinned at the bottom of CB3 for carrier densities larger than $\sim 5\times 10^{19}\text{cm}^{-3}$.

### v2

@@ Abstract
One convenient way to individually access the doping and energy levels of the surface and bulk electronic populations is to perform Landau level spectroscopy. The measurement of Shubnikov-de-Haas oscillations (SdHO) of the resistance under external magnetic field can give access to the carrier density, the transport and the quantum mobilities, the effective mass and the chemical potential of different charge carrier populations, whose contributions to the total conductance can be disentangled so. However, this technique requires magnetic fields such that $\mu B>1$, $\mu$ being the charge carrier mobility, different for each band. While such quantum oscillations have been studied in Sb-doped MnBi2Te4,[Jiang2021, Lee2021], they were not observed in pure MnBi2Te4 in the many magneto-transport studies up to 15 T[Cui2019, Lee2019, Chen2019a, Bac2022, Martini2023, Cao2024], except for the report of small resistance oscillations [Lei2022], which are not attributed to surface states, but to Weyl physics. This absence of SdHO in pure MnBi2Te4 is most probably due to a low mobility caused by the strong disorder in the bulk, attributed mainly to Mn/Bi antisites [Otrokov2019, Yan2019, Liu2020, Li2024b, Souchay2019, Zeugner2019], thus requiring very high magnetic fields to observe SdHO.
Figure 1: Device MBT2: Hall-bar geometry and metallic behavior. a: Atomic force microscopy picture of an etched MnBi2Te4 nanostructure, with thickness 91 nm and roughness of about 1 nm. The original shape of the MnBi2Te4 exfoliated flake can still be faintly seen below the contacts. b: temperature dependence of the longitudinal resistance in zero magnetic field. The antiferromagnetic transition induces a resistance peak at 25.5 K.
@@
[…]
Figure 2: Device MBT1 : Magneto-resistance and Hall response in the out-of-plane configuration. The longitudinal magnetoresistance (a) and the transverse magnetoresistance (b) were measured at 4.2K up to 55T, and the data were symmetrized and antisymmetrized with respect to the magnetic field, respectively. Shubnikov-de-Haas oscillations are visible above about 40 T. Purple dashed lines represent respectively a cubic fit of the magnetoresistance, used afterward to extract Shubnikov-de-Haas oscillations, and a linear fit of the Hall effect asymptote at high field. Arrows indicate the spin-flop- and saturation fields.
Figure 3: Evidence of Shubnikov-de-Haas oscillations and Lifshitz-Kosevich analysis. a: Residual Shubnikov-de-Haas oscillations after removal of a cubic background, at different temperatures. The position of the minimum and maximum of the oscillation are indicated by dashed lines. b: Residual Shubnikov-de-Haas oscillations after removal of a cubic background, versus inverse magnetic field. The dashed lines show the position of the SdHO extrema, separated by 0.003 T-1. c: Temperature dependence of the Shubnikov-de-Haas oscillation amplitude for the peaks at 43 and 49T. The error bars correspond to rapid noise in the raw data. Lifshitz-Kosevich fits lead effective masses $m^{*}$ of about $0.15-0.17m_{\mathrm{e}}$.
Fig.2 presents the longitudinal and the transverse magnetoresistance under perpendicular magnetic fields up to 55 T. At low field, both samples show the expected signature of a collinear antiferromagnet with a uniaxial anisotropy: a sharp resistance peak at about 3.1 T corresponds to the spin-flop magnetic transition,[Tan2020] which is also visible in the transverse resistance as a sharp jump of the anomalous Hall effect (blue arrows). Another feature is observed at 7 T, corresponding to the saturation field above which the magnetization is uniform (see Supplementary Information). This low-field magnetotransport, together with the observed Neel temperature, confirm that our nanostructures have all the usual magnetotransport properties expected of MnBi2Te4 samples. Above 10 T, the transverse resistance is linear up to 55 T, corresponding to an asymptote of the normal Hall effect, yielding a total bulk carrier density of $6.7\times 10^{19}$ cm-3.
@@
In order to determine the 3D or 2D character of the electronic population responsible for the observed SdHO, the magnetic field was tilted out of the sample plane by an angle $\theta$. From the angular dependence of the magnetoresistance we observe that the SdHO evolves with the angle, but is stable with respect to $B_{\perp}=B\cos(\theta)$. Fig.4a shows the SdHO plotted against the perpendicular component of the magnetic field $B_{\perp}$. The position of the maximum and minimum of the visible SdHO remains unaffected, while the positions vary with regard to the total field. This is a clear sign that the SdHO originate from a 2D electronic state in the sample plane. This is also compatible with the carrier density extracted from the SdHO frequency: if interpreted as a 3D parabolic bulk band, a magnetic frequency $f_{B}=167$ T corresponds to a bulk carrier density of $n_{\mathrm{3D}}^{\mathrm{SdHO}}=1/2\pi^{2}\,(2e/h\times f_{\mathrm{B}})^{3/2}=8\times 10^{17}$ cm-3, which is two orders of magnitude smaller than the total carrier density extracted from the Hall effect. Interpreting the SdHO as originating from a topological surface state (TSS) with linear energy dispersion and a Fermi velocity $v_{\mathrm{F}}=5.5\times 10^{5}$ m.s-1,[Tomarchio2022, Gong2019] we calculate a 2D carrier density of $n_{\mathrm{2D}}^{\mathrm{SdHO}}=e/h\times f_{\mathrm{B}}=4.1\times 10^{12}$ cm-2 at the topological surface state, and a Fermi energy $E_{\mathrm{F}}^{\mathrm{TSS}}=\hbar\cdot v_{\mathrm{F}}\cdot\sqrt{4\pi\cdot n_{\mathrm{2D}}^{\mathrm{SdHO}}}=$ 255 meV above the Dirac point in good agreement with ARPES [Xu2024, Estyunin2020].
Figure 4: Angular dependence of Shubnikov-de-Haas oscillations and band bending analysis. a: Angular dependence of the Shubnikov-de-Haas oscillations after removal of a cubic background, between 0°  (transverse field configuration) to 30°  tilt toward in-plane configuration, shown against the transverse magnetic field component $B_{\perp}=B\cdot\mathrm{cos}(\theta)$. Curves shifted for clarity. The schematic shows the tilting angle $\theta$ of the magnetic field. b: Schematic of the energy levels and band bending in the MnBi2Te4 nanostructure, as extracted from Shubnikov-de-Haas and Hall data (see main text), in a model with three conduction bands (CB1, CB2, CB3, see main text) and one topological surface state (TSS).
@@
We now model the chemical potential at the surface using the charge carrier density extracted from the SdH oscillations. As calculated above, the chemical potential of the surface lies 255 meV above the Dirac point. Using a separation of 100meV between the Dirac point and the lowest conduction band CB1 from the literature (as determined for instance by ARPES, see [Li2019f, Hao2019, Estyunin2020]), this imposes the position of the bulk band at the surface 155meV below the Fermi energy. As can be seen in Fig. 4, this results in an upward band bending of more than 150 meV between bulk and surface. This upward band bending is a common situation in heavily doped 3DTI [Veyrat2015], where topological surface states are filled by charge transfer from heavily doped bulk bands, resulting in a lower doping of the bulk band close to the topological surface over the length-scale of the Thomas-Fermi screening length. It is, however, interesting to notice that, given the important band bending between TSS and bulk states, surface-sensitive techniques such as ARPES measurements would significantly underestimate the bulk chemical potential, and in particular the filling of the third conduction band CB3. Given the large effective mass of CB3, the bulk chemical potential will efficiently be pinned at the bottom of CB3 for carrier densities larger than $\sim 5\times 10^{19}\text{cm}^{-3}$.
@@
Summing up the results of analysis and discussion, the hypothesis of 2D surface state-origin SdH oscillations is supported by:
a) the 1/cos angle dependence of the peak position in magnetic field,
b) the effective masses $m^{*}$ obtained from the model being consistent with the ones extracted from the Lifshitz-Kosevich fit of the measured SdH oscillations. This band bending model as obtained from the 3-band DFT model under the 2D hypothesis is further consistent with band bending reported on other topological insulators [Veyrat2015], and the surface chemical potential estimated from $f_{\mathrm{B}}$ is also consistent with ARPES results [Xu2024, Estyunin2020].